# Communication-Efficient Federated Fine-Tuning of Language Models via Dynamic Update Schedules

### v1

@@ 3.4. Default Configuration for FDA-Opt
(4) $\tilde{\tau}=2\cdot\tau+8\cdot\lceil e\rceil,\quad\text{where}\quad e\approx\frac{1}{|\mathcal{K}|}\sum_{k\in\mathcal{K}}|\mathcal{D}_{k}|$
Variance Query Indices. Within each round, we query the variance once per epoch. This completely alleviates the synchronization bottleneck encountered in the original work (Theologitis et al., 2025). We define:
@@ 4. EXPERIMENTAL APPROACH
4.1. Infrastructure
We conduct our experiments using the transformers library from Hugging Face (Wolf et al., 2020) with PyTorch (Paszke et al., 2019). The code for our implementation is available at https://github.com/miketheologitis/FDA-Opt. All experiments are performed on a local cluster equipped with 2 NVIDIA A10 GPUs and 46 Intel(R) Xeon(R) Silver 4310 CPUs.
4.2. Tasks & Datasets
Natural Language Understanding (NLU) tasks are a core component of evaluating machine learning models in NLP. These tasks test a model’s ability to understand, reason, and infer relationships between pieces of text. To conduct our experiments, we select six widely used datasets and tasks from the GLUE (Wang et al., 2018) benchmark: MRPC (Dolan and Brockett, 2005), SST-2 (Socher et al., 2013), RTE (Wang et al., 2018), QNLI (Wang et al., 2018), MNLI-m and MNLI-mm (Williams et al., 2018)—each corresponding to a distinct NLU task. Since the GLUE test sets are unpublished, we follow prior studies (Zhang et al., 2023) and use the validation sets as the new test sets.
4.3. Data Partitioning
@@
Clients. The scale of data federation is determined by the number of clients, which generally falls into two categories: cross-silo and cross-device (Kairouz et al., 2021; Qin et al., 2023). In cross-silo FL, the number of clients is small—such as hospitals or financial institutions—often allowing all clients to participate in each training round. For this setting, we use the MRPC and RTE datasets. In contrast, cross-device FL involves a much larger pool of clients, such as mobile or IoT devices, where only a fraction of them can participate in each round. For this setting, we use the SST-2, QNLI, and MNLI datasets. Table 2 summarizes the client partitioning and sampling strategies across these datasets.
Table 2. Client Partitioning Strategies per Dataset
@@
Label Distribution. One of the most common heterogeneity scenarios in FL is the presence of non-IID label distributions, where clients possess data with disproportionately distributed labels. For instance, one client may predominantly have samples of label A, while another may mostly contain label B. This phenomenon frequently occurs in real-world applications and is modeled using the Dirichlet distribution (Li et al., 2022; Lin et al., 2022; Qin et al., 2023). The degree of label imbalance is controlled by the concentration parameter $\alpha$. Larger values of $\alpha$ result in more uniform distributions, with $\alpha\to\infty$ producing a perfectly uniform distribution. Conversely, as $\alpha$ approaches zero, the distributions become highly imbalanced, with each client predominantly holding samples from a single label. In our experiments, we set $\alpha=1.0$, introducing a considerable level of heterogeneity.
@@
4.4. Models & Quality Baselines
In our experiments, we use the pre-trained transformer-based models RoBERTa (Liu et al., 2019) and DeBERTaV3 (He et al., 2023). More specifically, we use the roberta-base and microsoft/deberta-v3-base checkpoints available through Hugging Face (Wolf et al., 2020), which contain 125 million and 86 million parameters, respectively. Such encoder models are well-suited for FL, as they can be readily fine-tuned for a wide range of practical downstream NLP tasks (Lin et al., 2022) (e.g., sentiment analysis, textual entailment, etc.). We fine-tune both models with the FedOpt and FDA-Opt algorithms on the datasets outlined in Section 4.2.
Of course, our ultimate goal in FL is to train accurate and robust models. Thus, as in most of the FL literature (Reddi et al., 2021), we evaluate algorithms based on how efficiently they manage to do that. In other words, instead of comparing final accuracy scores or communication alone, we ask: how many communication rounds does it take to train a model with a specific target quality?
To define those targets in a consistent and task-agnostic way, we express them as percentages of the best-known centralized performance. Given both models’ widespread adoption, their top-reported performance when fine-tuned in centralized settings is well-documented (see Table 3). Naturally, due to the adverse conditions of data federation, FL performance is typically lower than centralized baselines. For example, the highest accuracy ever reported for RoBERTa when fine-tuned on MRPC is 90.2%; then, a target of 90% means attaining at least $81.18\%$ accuracy ($90\%\cdot 90.2=81.18\%$).
Table 3. Best-known accuracy scores ($\uparrow$) for fine-tuning each model on each task in an ideal centralized setting. These values serve as target performance baselines
RoBERTa (Liu et al., 2019) DeBERTa (He et al., 2023)
@@ 4.4. Models & Quality Baselines
4.5. Algorithms & Hyper-Parameters
We evaluate the most widely used algorithms from the FedOpt family and compare their performance with our proposed FDA-Opt algorithms. To ensure a fair, apples-to-apples comparison, each FedOpt algorithm is paired with its corresponding FDA-Opt variant using identical parameter configurations. For example, we compare FedAdamW with FDA-AdamW, keeping all optimizer settings fixed. The algorithmic pairs we compare are outlined in Table 1.
@@ 4.5. Algorithms & Hyper-Parameters
Lastly, unless otherwise specified, we set the number of local steps $\tau$ equal to one epoch—that is, $\tau=\lceil e\rceil$ from Equation (4)—as this is the most common and widely accepted choice in the FedOpt literature (Lin et al., 2022; Reddi et al., 2021; Hsu et al., 2019). Moreover, we use a batch size of $8$ and train for an exhaustive number of rounds, $T\in\{100,\dots,1000\}$, ensuring that training continues well beyond likely convergence.
5. EXPERIMENTAL RESULTS & ANALYSIS
@@ 5.1. Main Findings
5.2. Communication-Efficiency
Table 4. Comparison of FedOpt and FDA-Opt in communication-efficiency for training RoBERTa and DeBERTaV3 across different datasets to achieve target metrics. The table reports the number of rounds required to reach $85\%$, $90\%$, $95\%$, and $99\%$ of the top-reported results in the literature (see Table 3)—naturally, the fewer rounds, the better ($\downarrow$). For example, in the bottom-left cell, FedAdaGrad requires $29$ rounds to achieve $90\%$ of the best-reported accuracy for MRPC (i.e., at least $90\%\cdot 90.2=81.18\%$), while FDA-AdaGrad achieves the same in $21$ rounds, making it 1.4×1.4$\times$1.4 × more efficient. Similarly, it has a speedup of 1.5×1.5$\times$1.5 × for $95\%$. The “Avg. FDA Speedup” row is computed by averaging the individual FDA-Opt vs. FedOpt speedup values across each column
@@ 5.2. Communication-Efficiency
Avg. FDA Speedup 2.8×2.8$\times$2.8 × 2.3×2.3$\times$2.3 × 2.3×2.3$\times$2.3 × 2×2$\times$2 × 1.9×1.9$\times$1.9 × 2.3×2.3$\times$2.3 × 2.6×2.6$\times$2.6 × 1.6×1.6$\times$1.6 × 1.7×1.7$\times$1.7 × 1.6×1.6$\times$1.6 × 1.9×1.9$\times$1.9 × 1.6×1.6$\times$1.6 ×
@@
As is standard in the literature (Wang et al., 2021), we evaluate our algorithms by comparing their performance against predefined target metrics—such as accuracy ($\uparrow$). The primary objective of the algorithms is to train a model that attains the target metric. Once they first do, we assess their communication efficiency by examining the number of FL rounds they required ($\downarrow$). Table 4 illustrates this evaluation.
Cross-Silo. The cross-silo setting includes RoBERTa trained on the MRPC and RTE datasets, as outlined in Table 2.
  • RoBERTa on MRPC. As per Table 4, FDA-Opt demonstrates superior communication-efficiency, with an average speedup of 2.3×2.3$\times$2.3 ×-2.8×2.8$\times$2.8 × compared to FedOpt. It is 2.8×2.8$\times$2.8 × more efficient in reaching $90\%$ of the target metric, and 2.3×2.3$\times$2.3 × for $95\%$.
@@
Cross-Device. In the cross-device setting, we analyze performance of training RoBERTa on the SST-2 and QNLI datasets, and DeBERTaV3 on MNLI-m and MNLI-mm, as outlined in Table 2.
@@ 5.3. Convergence
Cross-Silo. On the MRPC dataset—which simulates a cross-silo setting following Table 2—all FDA-Opt algorithms demonstrate significantly faster convergence than their FedOpt counterparts (Figure 4). For instance, FDA-SGD achieves a training loss 100×100$\times$100 × smaller than FedAvg within the first $100$ rounds. Additionally, both FDA-SGDM and FDA-AdamW exhibit sharp drops within the initial 3030303040404040 rounds, outperforming FedAvgM and FedAdamW, respectively, with FDA-SGDM stabilizing at a loss 10×10$\times$10 × smaller than FedAvgM. Meanwhile, FDA-Adam converges to a loss 5×5$\times$5 × smaller than FedAdam, and FDA-AdaGrad consistently maintains a 5×5$\times$5 × lower loss than FedAdaGrad after the initial rounds.
Figure 4. Training Loss progression of training RoBERTa with FDA-Opt vs. FedOpt on MRPC (left) and SST-2 (right)
Cross-Device. On the SST-2 dataset—representing a cross-device setting as outlined in Table 2—all FDA-Opt algorithms consistently achieve 555510×10$\times$10 × lower training loss compared to their FedOpt counterparts (Figure 4). Unlike the cross-silo case, the initial drop in training loss is less pronounced, which is expected since each round involves only a small subset of clients. This leads to initial updates that are less representative of the overall training distribution.
5.4. Stability to Local Training Extension
Empirical evidence suggests that increasing the number of local training steps $\tau$ can lead to degraded performance or non-convergence altogether (Wang et al., 2021; Yu et al., 2019). Since our scheme extends local training even further ($\tilde{\tau}\gg\tau$), it is crucial to examine whether this extension introduces any instability. To this end, we train RoBERTa on MRPC (cross-silo) and SST-2 (cross-device) using a range of values for $\tau$.
@@ 5.4. Stability to Local Training Extension
Figure 5. Training RoBERTa with FDA-Opt vs. FedOpt for varying initial local training steps $\tau$ (measured in epochs). We report the number of rounds required to attain the target accuracy. Failure to converge is marked with “$\bm{\times}$
6. RELATED WORK
@@
Round Termination. Several works have explored strategies for determining the round termination intervals in FL. The study in (Wang et al., 2019) aims to find a fixed, optimal interval by balancing local computation and aggregation constraints. In contrast, (Mills et al., 2023) analyzes the trade-off between fast convergence and round completion rates under strongly convex client objectives, leading to a decaying round duration scheme. Similarly, (Wang and Joshi, [n. d.]) minimizes convergence error with respect to wall-clock time by progressively decreasing the round duration. On the other hand, (Haddadpour et al., 2019) focuses on minimizing communication rounds for a fixed number of model updates, which results in an increasing sequence of round duration intervals. Unlike these approaches, which rely on predefined schedules—whether fixed, decaying, or increasing—FDA-Opt makes round termination decisions dynamically, in real-time, based on the state of the training.
@@
Convergence. The most direct way to alleviate the communication burden in FL is to accelerate convergence. This motivated the development of the FedOpt family (Section 2.2). However, FL algorithms still struggle with high heterogeneity, where different clients converge to disparate and often conflicting local minima (Karimireddy et al., 2020; Reddi et al., 2021; Wenig and Papenbrock, 2022), ultimately slowing down convergence. To address this, SCAFFOLD introduced control variates to correct client drift (Karimireddy et al., 2020). Similarly, the Mime framework (Karimireddy et al., 2021) leverages both control variates and server statistics. FedProx(Li et al., 2020) tackles heterogeneity by adding an $L^{2}$ regularization term. Notably, it has been demonstrated that FedOpt outperforms FedProx in NLP tasks similar to ours (Lin et al., 2022). Lastly, FedDyn (Acar et al., 2021) introduces a dynamic regularizer, ensuring that client convergence aligns with the stationary point of the global objective. Importantly, FDA-Opt is orthogonal to these optimization methods as we focus on a complementary aspect: when to terminate rounds.
@@
PEFT. Parameter-efficient fine-tuning (PEFT) (Li and Liang, 2021; Lester et al., 2021; Hu et al., 2022; Ben Zaken et al., 2022) trains only a small subset of a model’s parameters while keeping the rest frozen. One of the most widely used approaches, LoRA (Hu et al., 2022), injects low-rank adaptation matrices into a transformer’s frozen attention layers, reducing the number of trainable parameters to $<1\%$. Given the communication constraints in FL, recent studies have explored the applicability of PEFT methods in this setting (Zhang et al., 2023; Sun et al., 2024). However, techniques like LoRA struggle with the high heterogeneity present in FL (Sun et al., 2024). SLoRA (Babakniya et al., 2023) addresses this issue by leveraging sparse fine-tuning (Ansell et al., 2022) to initialize the injected parameters more effectively. These PEFT methods can be readily integrated with FDA-Opt, as the choice of trainable parameters, $\mathbf{w}$, remains user-defined.
@@
Compression. Another way to reduce communication in FL is to compress transmitted information using techniques like quantization (Shlezinger et al., 2021) and sparsification (Aji and Heafield, 2017; Basu et al., 2019). A comprehensive survey of these methods is provided in (Wang et al., 2023). Notably, compression is orthogonal to our approach and can be seamlessly integrated with FDA-Opt.
7. CONCLUSION

### v2

@@ III-D Default Configuration for FDA-Opt
$\tilde{\tau}=2\cdot\tau+8\cdot\lceil e\rceil,\quad\text{where}\quad e\approx\frac{1}{|\mathcal{K}|}\sum_{k\in\mathcal{K}}|\mathcal{D}_{k}|$ (4)
Variance Query Indices. Within each round, we query the variance once per epoch. This completely alleviates the synchronization bottleneck encountered in the original work [13]. We define:
@@ IV Experimental Approach
IV-A Infrastructure
We conduct our experiments using the transformers library from Hugging Face [20] with PyTorch [21]. The code is available at GitHub [^2] https://github.com/miketheologitis/FDA-Opt. All experiments are performed on a local cluster equipped with 2 NVIDIA A10 GPUs and 46 Intel(R) Xeon(R) Silver 4310 CPUs.
IV-B Tasks & Datasets
Natural Language Understanding (NLU) tasks are a core component of evaluating machine learning models in NLP. These tasks test a model’s ability to understand, reason, and infer relationships between pieces of text. To conduct our experiments, we select six widely used datasets and tasks from the GLUE [22] benchmark: MRPC [23], SST-2 [24], RTE [22], QNLI [22], MNLI-m and MNLI-mm [25]—each corresponding to a distinct NLU task. Since the GLUE test sets are unpublished, we follow prior studies [26] and use the validation sets as the new test sets.
IV-C Data Partitioning
@@
Clients. The scale of data federation is determined by the number of clients, which generally falls into two categories: cross-silo and cross-device [27]. In cross-silo FL, the number of clients is small—such as hospitals or financial institutions—often allowing all clients to participate in each training round. For this setting, we use the MRPC and RTE datasets. In contrast, cross-device FL involves a much larger pool of clients, such as mobile or IoT devices, where only a fraction of them can participate in each round. For this setting, we use the SST-2, QNLI, and MNLI datasets. Table II summarizes the client partitioning and sampling strategies across these datasets.
TABLE II: Client Partitioning Strategies per Dataset
@@
Label Distribution. One of the most common heterogeneity scenarios in FL is the presence of non-IID label distributions, where clients possess data with disproportionately distributed labels. For instance, one client may predominantly have samples of label A, while another may mostly contain label B. This phenomenon frequently occurs in real-world applications and is modeled using the Dirichlet distribution [28, 29, 30]. The degree of label imbalance is controlled by the concentration parameter $\alpha$. Larger values of $\alpha$ result in more uniform distributions, with $\alpha\to\infty$ producing a perfectly uniform distribution. Conversely, as $\alpha$ approaches zero, the distributions become highly imbalanced, with each client predominantly holding samples from a single label. In our experiments, we set $\alpha=1.0$, introducing a considerable level of heterogeneity.
@@
IV-D Models & Quality Baselines
In our experiments, we use the pre-trained transformer-based models RoBERTa [31] and DeBERTaV3 [32]. More specifically, we use the roberta-base and microsoft/deberta-v3-base checkpoints available through Hugging Face [20], which contain 125 million and 86 million parameters, respectively. Such encoder models are well-suited for FL, as they can be readily fine-tuned for a wide range of practical downstream NLP tasks [29] (e.g., sentiment analysis, textual entailment, etc.). We fine-tune both models with the FedOpt and FDA-Opt algorithms on the datasets outlined in Section IV-B.
Of course, our ultimate goal in FL is to train accurate and robust models. Thus, as in most of the FL literature [11], we evaluate algorithms based on how efficiently they manage to do that. In other words, instead of comparing final accuracy scores or communication alone, we ask: how many communication rounds does it take to train a model with a specific target quality?
To define those targets in a consistent and task-agnostic way, we express them as percentages of the best-known centralized performance. Given both models’ widespread adoption, their top-reported performance when fine-tuned in centralized settings is well-documented (see Table III). Naturally, due to the adverse conditions of data federation, FL performance is typically lower than centralized baselines. For example, the highest accuracy ever reported for RoBERTa when fine-tuned on MRPC is 90.2%; then, a target of 90% means attaining at least $81.18\%$ accuracy ($90\%\cdot 90.2=81.18\%$).
TABLE III: Best-known accuracy scores ($\uparrow$) for fine-tuning each model on each task in an ideal centralized setting. These values serve as target performance baselines
RoBERTa [31] DeBERTa [32]
@@ IV-D Models & Quality Baselines
IV-E Algorithms & Hyper-Parameters
We evaluate the most widely used algorithms from the FedOpt family and compare their performance with our proposed FDA-Opt algorithms. To ensure a fair, apples-to-apples comparison, each FedOpt algorithm is paired with its corresponding FDA-Opt variant using identical parameter configurations. For example, we compare FedAdamW with FDA-AdamW, keeping all optimizer settings fixed. The algorithmic pairs we compare are outlined in Table I.
@@ IV-E Algorithms & Hyper-Parameters
Lastly, unless otherwise specified, we set the number of local steps $\tau$ equal to one epoch—that is, $\tau=\lceil e\rceil$ from Equation (4)—as this is the most common and widely accepted choice in the FedOpt literature [29, 11, 15]. Moreover, we use a batch size of $8$ and train for an exhaustive number of rounds, $T\in\{100,\dots,1000\}$, ensuring that training continues well beyond likely convergence.
V Experimental Results & Analysis
@@ V-A Main Findings
V-B Communication-Efficiency
TABLE IV: Comparison of FedOpt and FDA-Opt in communication-efficiency for training RoBERTa and DeBERTaV3 across different datasets to achieve target metrics. The table reports the number of rounds required to reach $85\%$, $90\%$, $95\%$, and $99\%$ of the top-reported results in the literature (see Table III)—naturally, the fewer rounds, the better ($\downarrow$). For example, in the bottom-left cell, FedAdaGrad requires $29$ rounds to achieve $90\%$ of the best-reported accuracy for MRPC (i.e., at least $90\%\cdot 90.2=81.18\%$), while FDA-AdaGrad achieves the same in $21$ rounds, making it 1.4×1.4$\times$1.4 × more efficient. Similarly, it has a speedup of 1.5×1.5$\times$1.5 × for $95\%$. The “Avg. FDA-Opt Speedup” row is computed by averaging the individual FDA-Opt vs. FedOpt speedup values across each column
@@ V-B Communication-Efficiency
Avg. FDA-Opt Speedup $\bm{2.8\times}$ $\bm{2.3\times}$ $\bm{2.3\times}$ $\bm{2\times}$ $\bm{1.9\times}$ $\bm{2.3\times}$ $\bm{2.6\times}$ $\bm{1.6\times}$ $\bm{1.7\times}$ $\bm{1.6\times}$ $\bm{1.9\times}$ $\bm{1.6\times}$
@@
As is standard in the literature [33], we evaluate our algorithms by comparing their performance against predefined target metrics—such as accuracy ($\uparrow$). The primary objective of the algorithms is to train a model that attains the target metric. Once they first do, we assess their communication efficiency by examining the number of FL rounds they required ($\downarrow$). Table IV illustrates this evaluation.
Cross-Silo. The cross-silo setting includes RoBERTa trained on the MRPC and RTE datasets, as outlined in Table II.
  • RoBERTa on MRPC. As per Table IV, FDA-Opt demonstrates superior communication-efficiency, with an average speedup of 2.3×2.3$\times$2.3 ×-2.8×2.8$\times$2.8 × compared to FedOpt. It is 2.8×2.8$\times$2.8 × more efficient in reaching $90\%$ of the target metric, and 2.3×2.3$\times$2.3 × for $95\%$.
@@
Cross-Device. In the cross-device setting, we analyze performance of training RoBERTa on the SST-2 and QNLI datasets, and DeBERTaV3 on MNLI-m and MNLI-mm, as outlined in Table II.
@@ V-C Convergence
Cross-Silo. On the MRPC dataset—which simulates a cross-silo setting following Table II—all FDA-Opt algorithms demonstrate significantly faster convergence than their FedOpt counterparts (Figure 4). For instance, FDA-SGD achieves a training loss 100×100$\times$100 × smaller than FedAvg within the first $100$ rounds. Additionally, both FDA-SGDM and FDA-AdamW exhibit sharp drops within the initial 3030303040404040 rounds, outperforming FedAvgM and FedAdamW, respectively, with FDA-SGDM stabilizing at a loss 10×10$\times$10 × smaller than FedAvgM. Meanwhile, FDA-Adam converges to a loss 5×5$\times$5 × smaller than FedAdam, and FDA-AdaGrad consistently maintains a 5×5$\times$5 × lower loss than FedAdaGrad after the initial rounds.
Figure 4: Training Loss progression of training RoBERTa with FDA-Opt vs. FedOpt on MRPC (left) and SST-2 (right)
Cross-Device. On the SST-2 dataset—representing a cross-device setting as outlined in Table II—all FDA-Opt algorithms consistently achieve 555510×10$\times$10 × lower training loss compared to their FedOpt counterparts (Figure 4). Unlike the cross-silo case, the initial drop in training loss is less pronounced, which is expected since each round involves only a small subset of clients. This leads to initial updates that are less representative of the overall training distribution.
V-D Stability to Local Training Extension
Empirical evidence suggests that increasing the number of local training steps $\tau$ can lead to degraded performance or non-convergence altogether [33, 12]. Since our scheme extends local training even further ($\tilde{\tau}\gg\tau$), it is crucial to examine whether this extension introduces any instability. To this end, we train RoBERTa on MRPC (cross-silo) and SST-2 (cross-device) using a range of values for $\tau$.
@@ V-D Stability to Local Training Extension
Figure 5: Training RoBERTa with FDA-Opt vs. FedOpt for varying initial local training steps $\tau$ (measured in epochs). We report the number of rounds required to attain the target accuracy. Failure to converge is marked with “$\bm{\times}$
VI Related Work
@@
Round Termination. Several works have explored strategies for determining the round termination intervals in FL. The study in [34] aims to find a fixed, optimal interval by balancing local computation and aggregation constraints. In contrast, [35] analyzes the trade-off between fast convergence and round completion rates under strongly convex client objectives, leading to a decaying round duration scheme. Similarly, [36] minimizes convergence error with respect to wall-clock time by progressively decreasing the round duration. On the other hand, [37] focuses on minimizing communication rounds for a fixed number of model updates, which results in an increasing sequence of round duration intervals. Unlike these approaches, which rely on predefined schedules—whether fixed, decaying, or increasing—FDA-Opt makes round termination decisions dynamically, in real-time, based on the state of the training.
@@
Convergence. The most direct way to alleviate the communication burden in FL is to accelerate convergence. This motivated the development of the FedOpt family (Section II-B). However, FL algorithms still struggle with high heterogeneity, where different clients converge to disparate and often conflicting local minima [38, 11, 39], ultimately slowing down convergence. To address this, SCAFFOLD introduced control variates to correct client drift [38]. Similarly, the Mime framework [40] leverages both control variates and server statistics. FedProx[41] tackles heterogeneity by adding an $L^{2}$ regularization term. Notably, it has been demonstrated that FedOpt outperforms FedProx in NLP tasks similar to ours [29]. Lastly, FedDyn [42] introduces a dynamic regularizer, ensuring that client convergence aligns with the stationary point of the global objective. Importantly, FDA-Opt is orthogonal to these optimization methods as we focus on a complementary aspect: when to terminate rounds.
@@
PEFT. Parameter-efficient fine-tuning (PEFT) [43, 44, 45, 46] trains only a small subset of a model’s parameters while keeping the rest frozen. One of the most widely used approaches, LoRA [45], injects low-rank adaptation matrices into a transformer’s frozen attention layers, reducing the number of trainable parameters to $<1\%$. Given the communication constraints in FL, recent studies have explored the applicability of PEFT methods in this setting [26, 47]. However, techniques like LoRA struggle with the high heterogeneity present in FL [47]. SLoRA [48] addresses this issue by leveraging sparse fine-tuning [49] to initialize the injected parameters more effectively. These PEFT methods can be readily integrated with FDA-Opt, as the choice of trainable parameters, $\mathbf{w}$, remains user-defined.
@@
Compression. Another way to reduce communication in FL is to compress transmitted information using techniques like quantization [50] and sparsification [51, 52]. A comprehensive survey of these methods is provided in [53]. Notably, compression is orthogonal to our approach and can be seamlessly integrated with FDA-Opt.
VII Conclusion